Solve each system by graphing.
y=3x+6 and Y=1/5x-8

Answer by Alan3354(69443)

Solve each system by graphing.
y=3x+6 and Y=1/5x-8
==============
Is the 2nd eqn y = (1/5)x - 8 ?
Or y = 1/(5x) - 8 ?
Or y = 1/(5x - 8) ?
========================
Parentheses are free. Use some.